40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 27 16:26:11 2019, Yehonathan, Update, PSL, PMC Linewidth measurement PMCTransmissionSpectra.pdf
    Reply  Wed Jan 1 14:14:00 2020, Yehonathan, Update, PSL, PMC Linewidth measurement 150MHzSideBandCreation.pdfLinewidthMeasurment.pdf
       Reply  Wed Jan 15 15:04:31 2020, gautam, Update, PSL, PMC Linewidth measurement 
          Reply  Fri Jan 17 22:11:19 2020, Yehonathan, Update, PSL, Ringdown measurements 20200117_192455.jpg20200117_192448.jpg
             Reply  Sat Jan 25 11:54:42 2020, Yehonathan, Update, PSL, Ringdown measurements Input_beam_path.jpgReflected_Beam_Path.jpgTransmitted_Beam_Path.jpgPMCRingdownNormalizedRawdata.pdfTransPDFits.pdf
                Reply  Mon Jan 27 21:35:06 2020, Yehonathan, Update, PSL, Ringdown measurements IMC_REFL_Beam_Path.jpg
                   Reply  Mon Jan 27 21:48:49 2020, gautam, Update, PSL, Ringdown measurements elevatedSeis.pdf
                      Reply  Tue Jan 28 16:01:17 2020, gautam, Update, IOO, IMC WFS servos stable again QPD_recenter.png
                         Reply  Tue Jan 28 20:51:37 2020, Yehonathan, Update, IOO, IMC WFS servos stable again 
                            Reply  Wed Jan 29 12:40:24 2020, Yehonathan, Update, IOO, IMC Ringdowns preliminary data analysis IMCRingdownNormalizedRawdata.pdfIMCTransPDFits.pdf
                               Reply  Mon Feb 3 13:54:10 2020, Yehonathan, Update, IOO, IMC Ringdowns extended data analysis IMCTransReflAnalysis_anotated.pdf
                                  Reply  Wed Feb 5 21:13:17 2020, Yehonathan, Update, IOO, IMC Ringdowns extended data analysis 
                                     Reply  Sat Feb 15 12:56:24 2020, Yehonathan, Update, IOO, IMC Transfer function measurement IMCFrequencyResponse.pdfIMCRingdownNormalizedRawdata.pdfIMCREFLPDFits.pdf
Message ID: 15165     Entry time: Tue Jan 28 16:01:17 2020     In reply to: 15161     Reply to this: 15170
Author: gautam 
Type: Update 
Category: IOO 
Subject: IMC WFS servos stable again 

With all of the shaking (man-made and divine), it was a hard to debug this problem. Summary of fixes:

  1. The beam was misaligned on the WFS 1 and 2 heads, as well as the MC2 trans QPD. I re-aligned the former with the IMC unlocked, the latter (see Attachment) with the IMC locked (but the MC2 spot centering loops disabled).
  2. I reset the WFS DC and RF offsets, as well as the QPD offsets (once I had hand-aligned the IMC mirrors to obtain good transmission).

At least the DC indicators are telling me that the IMC locking is back to a somewhat stable state. I have not yet checked the frequency noise / RIN.

Attachment 1: QPD_recenter.png  53 kB  | Hide | Hide all
QPD_recenter.png
ELOG V3.1.3-